# Exact Thresholds for Noisy Non-Adaptive Group Testing

### 10 Jan 2024 v1

@@ I.2 Near-Constant Weight Design
Recall that each item is placed in $\Delta:=\frac{\nu n}{k}$ tests for some $\nu=\Theta(1)$, and that the information density $\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})$ is defined in (18). Throughout much of the analysis in this appendix, it will be useful to condition on $\mathbf{X}_{s_{\mathrm{eq}}}$ (or its specific realization $\mathbf{x}_{s_{\mathrm{eq}}}$ depending on the context), and we accordingly adopt the shorthand $\mathsf{P}(\cdot)=\mathbb{P}(\cdot|\mathbf{X}_{s_{\mathrm{eq}}})$ (or $\mathsf{P}(\cdot)=\mathbb{P}(\cdot|\mathbf{x}_{s_{\mathrm{eq}}})$).
Given $\mathbf{X}_{s_{\mathrm{eq}}}$, we divide $\mathbf{Y}$ into $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$, where $\mathbf{Y}_{1}$ represents the results of tests corresponding to $X_{s_{\mathrm{eq}}}^{(i)}\neq 0$, and $\mathbf{Y}_{2}$ represents the remaining results corresponding to $X_{s_{\mathrm{eq}}}^{(i)}=0$. Due to the “OR” operation in the group testing model, $\mathbf{Y}_{1}$ is completely determined by the noise in its tests, which implies that $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ are conditionally independent given $\mathbf{X}_{s_{\mathrm{eq}}}$. Therefore, we have
$\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})$ $=\log\frac{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})}{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})}$ (352)
$=\log\frac{\mathsf{P}(\mathbf{Y}_{1}|\mathbf{X}_{s_{\mathrm{dif}}})\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})}{\mathsf{P}(\mathbf{Y}_{1})\mathsf{P}(\mathbf{Y}_{2})}$ (353)
$=\log\frac{\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})}{\mathsf{P}(\mathbf{Y}_{2})},$ (354)
where (354) follows from $\mathsf{P}(\mathbf{Y}_{1}|\mathbf{X}_{s_{\mathrm{dif}}})=\mathsf{P}(\mathbf{Y}_{1})$ due to the fact that $\mathbf{Y}_{1}$ only depends on the noise. We let $n_{i}$ be the length of $\mathbf{Y}_{i}$ ($i=1,2$), which is deteriministic given $\mathbf{X}_{s_{\mathrm{eq}}}$.
J.1 Proof of Lemma 11 (Concentration Bound for $\frac{\ell}{k}\to\alpha\in[0,1]$)
@@
Concentration behavior of various quantities: We first consider the case $\alpha\in(0,1]$ and establish (349). Recall that we let $n_{i}$ be the length of $\mathbf{Y}_{i}$ for $i=1,2$. Let $M$ denote the number of tests corresponding to “$X_{s_{\mathrm{eq}}}^{(i)}=0,X_{s_{\mathrm{dif}}}^{(i)}\neq 0$”, [^10] Here, $M$ represents the random variable, while we will use lowercase letter $m$ for its specific value or actual realization; we will use similar convention for other variables below, such as $V$ and $v$. meaning that $n_{2}-M$ is the number of tests corresponding to “$X_{s_{\mathrm{eq}}}^{(i)}=0,X_{s_{\mathrm{dif}}}^{(i)}=0$”, or equivalently $X^{(i)}_{s}=0$. This definition of $M$ coincides with $M_{s_{\mathrm{dif}}}$ in the notation introduced above Lemma 3, but we omit the subscript to lighten notation, since we only consider one particular choice of $s_{\mathrm{dif}}$ here.
Note that $\frac{|s_{\mathrm{eq}}|}{k}=\frac{k-\ell}{k}=1-\frac{\ell}{k}$, and $\mathbf{Y}_{1}$ corresponds to the tests where “$X_{s_{\mathrm{eq}}}^{(i)}\neq 0$” so that we can write $n_{1}=W^{(s_{\mathrm{eq}})}$ in the notation of Lemma 7. Thus, for some $\delta>0$, (319) and (321) in Lemma 7 give that with probability at least $1-2\exp\big{(}-2\nu^{-1}\delta^{2}n\big{)}$, we have
$n_{1}$ $=\Big{(}1-e^{-\nu(1-\frac{\ell}{k})(1+o(1))}+\delta^{\prime}\sqrt{1-\frac{\ell}{k}}\Big{)}n,\text{ for some }|\delta^{\prime}|<\delta,$ (355)
@@
$n_{2}=n-n_{1}=\Big{(}e^{-\nu(1-\frac{\ell}{k})(1+o(1))}-\delta^{\prime}\sqrt{1-\frac{\ell}{k}}\Big{)}n,\text{ for some }|\delta^{\prime}|<\delta.$ (356)
We note that (356) holds for all $\ell=1,\dotsc,k$. For the case of $\frac{\ell}{k}\to\alpha\in(0,1]$, we obtain from (356) that the event
$n_{2}=\big{(}e^{-\nu(1-\alpha)}+o(1)-\delta^{\prime}\big{)}n,\text{ for some }|\delta^{\prime}|<\delta.$ (357)
@@
Next, we seek to obtain a similar concentration result for $M$. Note that $n-(n_{2}-M)$ represents the number of tests where $X_{s}^{(i)}\neq 0$, so we can write $n-(n_{2}-M)=W^{(s)}$. Hence, (319) and (321) in Lemma 7 yield that with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$, we have
$n-(n_{2}-M)=(1-e^{-\nu}+o(1)+\delta^{\prime\prime})n,\leavevmode\nobreak\ \text{for some }|\delta^{\prime\prime}|<\delta.$ (358)
Then, (357) and (358) imply
$M=(e^{-\nu(1-\alpha)}-e^{-\nu}+o(1)+\delta^{(3)})n,\leavevmode\nobreak\ \text{for some }|\delta^{(3)}|<2\delta.$ (359)
In the subsequent analysis, we will frequently suppose that (357) and (359) hold, which is justified via a simple union bound for the multiple high-probability events.
@@
We take $t=n_{2}\delta$ to obtain that $|V-[M(1-\rho)+(n_{2}-M)\rho]|<n_{2}\delta$, or equivalently, $V=(1-2\rho)M+(\rho+\delta^{(4)})n_{2}$ for some $|\delta^{(4)}|<\delta$, with probability at least $1-2\exp\big{(}-2n_{2}\delta^{2}\big{)}$. Substituting the high-probability values of $M$ and $n_{2}$ in (357) and (359), when $\delta<\frac{1}{4}e^{-\nu(1-\alpha)}$, with probability at least $1-2\exp(-e^{-\nu(1-\alpha)}\delta^{2}n)$ we have
$V$ $=(1-2\rho)\big{(}e^{-\nu(1-\alpha)}-e^{-\nu}+o(1)+\delta^{(3)}\big{)}n+(\rho+\delta^{(4)})\big{(}e^{-\nu(1-\alpha)}+o(1)-\delta^{\prime}\big{)}n$ (361)
$=ne^{-\nu(1-\alpha)}\Big{(}(1-2\rho)(1-e^{-\nu\alpha})+\rho+o(1)+c_{1,\rho,\nu,\alpha}\delta\Big{)}$ (362)
$=ne^{-\nu(1-\alpha)}\big{(}\rho\star e^{-\nu\alpha}+o(1)+c_{1,\rho,\nu,\alpha}\delta\big{)},$ (363)
@@
where in (362) we introduce $c_{1,\rho,\nu,\alpha}$ that is bounded by an absolute constant depending on $(\rho,\nu,\alpha)$, [^11] Similarly, we will use $c_{2,\rho,\nu,\alpha}$, $c_{3,\rho,\nu,\alpha}$, etc. to denote constants depending on $(\rho,\nu,\alpha)$. in (363) we make the observation $(1-2\rho)(1-e^{-\nu\alpha})+\rho=\rho\star e^{-\nu\alpha}$ (recalling that $a\star b=ab+(1-a)(1-b)$).
@@
Recall that (354) gives $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})=\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})-\log\mathsf{P}(\mathbf{Y}_{2})$, and in both terms we implicitly condition on $\mathbf{X}_{s_{\mathrm{eq}}}$. To establish the concentration bound for $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})$, the main work is to bound these two terms separately.
@@
Bounding $\log\mathsf{P}(\mathbf{Y}_{2})$: To handle the term $\mathsf{P}(\mathbf{Y}_{2})$, consider a fixed realization $\mathbf{y}_{2}$ of $\mathbf{Y}_{2}$. This implies that the realization $v$ of $V$ is also fixed (since it is a function of $\mathbf{y}_{2}$), and we suppose that it satisfies the high-probability condition (363). By the symmetry of the test design with respect to re-ordering the test indices, all $\binom{n_{2}}{v}$ possibilities of $\mathbf{Y}_{2}$ having weight $v$ occur with equal probability, and hence, we have $\mathsf{P}(\mathbf{y}_{2})=\frac{\mathsf{P}(v)}{\binom{n_{2}}{v}}$. Then, applying [^12] This step may appear loose, but the idea is that if we were to consider all $v$ and take the dominant one, it would have a “large” value of $\mathbb{P}(v)$ anyway, i.e., not decaying fast enough to impact the final result. $\mathsf{P}(v)\leq 1$ and substituting (357) and (363), we obtain
$\mathsf{P}(\mathbf{y}_{2})\leq\frac{1}{\binom{ne^{-\nu(1-\alpha)}(1+o(1)+c_{2,\rho,\nu,\alpha}\delta)}{ne^{-\nu(1-\alpha)}(e^{-\alpha\nu}\star\rho+o(1)+c_{1,\rho,\nu,\alpha}\delta)}}.$ (364)
@@
From Lemma 2, we have $\binom{N}{\beta N}\geq\frac{\exp(NH_{2}(\beta))}{\sqrt{2N}}=\exp(N[H_{2}(\beta)+o(1)])$ for constant $\beta\in(0,1)$, which implies that if $\delta$ is smaller than some constant depending on $(\alpha,\nu,\rho)$ such that $\frac{e^{-\alpha\nu}\star\rho+o(1)+c_{1,\rho,\nu,\alpha}\delta}{1+o(1)+c_{2,\rho,\nu,\alpha}\delta}$ is sufficiently close to $e^{-\alpha\nu}\star\rho$, we can upper bound the right-hand side of (364) by $\exp\big{(}-[1+o(1)+c_{3,\rho,\nu,\alpha}\delta]ne^{-\nu(1-\alpha)}H_{2}(e^{-\alpha\nu}\star\rho)\big{)}$ for some constant $c_{3,\rho,\nu,\alpha}$. Since this holds for any realization of $\mathbf{y}_{2}$ whose corresponding $v$ value satisfies (363), we conclude that
$\mathsf{P}(\mathbf{Y}_{2})\leq\exp\Big{(}-[1+o(1)+c_{3,\rho,\nu,\alpha}\delta]ne^{-\nu(1-\alpha)}H_{2}(e^{-\alpha\nu}\star\rho)\Big{)}$ (365)
@@
$\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})=\log\prod_{i=1}^{n_{2}}\mathsf{P}(Y_{2}^{(i)}|X_{s_{\mathrm{dif}}}^{(i)})=\sum_{i=1}^{n_{2}}\log\mathsf{P}(Y_{2}^{(i)}|X_{s_{\mathrm{dif}}}^{(i)}),$ (366)
where without loss of generality we index the tests in $\mathbf{Y}_{2}$ from $1$ to $n_{2}$. Note that (366) indicates that $\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})$ is independent sum of $n_{2}$ random variables, which have a common expected value $-H_{2}(\rho)$ and are bounded by $|\log\rho|$ (since for any $\rho\in\big{(}0,\frac{1}{2}\big{)}$ we have $|\rho\log\rho+(1-\rho)\log(1-\rho)|\leq\rho|\log\rho|+(1-\rho)|\log\rho|=|\log\rho|$), so using Hoeffding’s inequality [12, Thm. 2.8] with respect to the noise randomness gives the following for any $t>0$:
$\mathsf{P}\Big{(}\big{|}\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})+n_{2}H_{2}(\rho)\big{|}\geq t\Big{)}\leq 2\exp\Big{(}-\frac{2t^{2}}{n_{2}\log^{2}\rho}\Big{)}.$ (367)
Moreover, we use (367) and let $t=\delta n_{2}H_{2}(\rho)$ to obtain that $|\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})+n_{2}H_{2}(\rho)|\leq\delta n_{2}H_{2}(\rho)$ with probability at least $1-2\exp\big{(}-2\delta^{2}(\log\rho)^{-2}(H_{2}(\rho))^{2}n_{2}\big{)}$. Further substituting (357), when $\delta<\frac{1}{4}e^{-\nu(1-\alpha)}$, we obtain that
$\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})=-ne^{-\nu(1-\alpha)}H_{2}(\rho)\big{(}1+c_{4,\rho,\nu,\alpha}\delta\big{)}$ (368)
@@
Putting the pieces together: We have shown that (365) and (368) hold with probability at least $1-\exp\big{(}-C_{\rho,\nu,\alpha}\delta^{2}n\big{)}$ for some $C_{\rho,\nu,\alpha}>0$ depending on $(\rho,\nu,\alpha)$ only (here we suppose that $\delta$ is given and $n$ is large enough, so that any leading constant before $\exp(-C_{\rho,\nu,\alpha}\delta^{2}n)$ can be absorbed into the exponent). Now we start with (354) and proceed as
$\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})=\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})-\log\mathsf{P}(\mathbf{Y}_{2})$ (369)
$\geq-ne^{-\nu(1-\alpha)}H_{2}(\rho)(1+c_{4,\rho,\nu,\alpha}\delta)+[1+o(1)+c_{3,\rho,\nu,\alpha}\delta]ne^{-\nu(1-\alpha)}H_{2}(e^{-\alpha\nu}\star\rho)$ (370)
$\geq(1-o(1))I_{\ell}^{n}+\big{(}c_{3,\rho,\nu,\alpha}H_{2}(e^{-\alpha\nu}\star\rho)-c_{4,\rho,\nu,\alpha}H_{2}(\rho)\big{)}\delta ne^{-\nu(1-\alpha)}$ (371)
$\geq(1-o(1))I_{\ell}^{n}-C_{1,\rho,\nu,\alpha}n\delta$ (372)
$=\Big{(}1-o(1)-\frac{C_{1,\rho,\nu,\alpha}n\delta}{I_{\ell}^{n}}\Big{)}I_{\ell}^{n},$ (373)
where in (370) we substitute (365) and (368), in (371) we recall $I_{\ell}^{n}$ given in (348) in Lemma 10, and in (372) $C_{1,\rho,\nu,\alpha}$ is a constant depending on $(\rho,\nu,\alpha)$. Thus, given any $\delta_{1}>0$, we can set the above $\delta$ as $\delta=\frac{I_{\ell}^{n}\delta_{1}}{2nC_{1,\rho,\nu,\alpha}}$, and we obtain $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})\geq(1-o(1)-\frac{\delta_{1}}{2})I_{\ell}^{n}$ with probability at least $1-4\exp\big{(}-\frac{C_{\rho,\nu,\alpha}^{\prime}(I_{\ell}^{n})^{2}\delta_{1}^{2}}{n}\big{)}$ for some $C^{\prime}_{\rho,\nu,\alpha}>0$ only depending on $(\rho,\nu,\alpha)$, as desired.

### 23 Oct 2024 v2

@@ I.2 Near-Constant Weight Design
Recall that each item is placed in $\Delta:=\frac{\nu n}{k}$ tests for some $\nu=\Theta(1)$, and that the information density $\imath^{n}(\mathbf{X}_{s_{\mathrm{dif}}};\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})$ is defined in (3.1). Throughout much of the analysis in this appendix, it will be useful to condition on $\mathbf{X}_{s_{\mathrm{eq}}}$ (or its specific realization $\mathbf{x}_{s_{\mathrm{eq}}}$ depending on the context), and we accordingly adopt the shorthand $\mathsf{P}(\cdot)=\mathbb{P}(\cdot|\mathbf{X}_{s_{\mathrm{eq}}})$ (or $\mathsf{P}(\cdot)=\mathbb{P}(\cdot|\mathbf{x}_{s_{\mathrm{eq}}})$).
Given $\mathbf{X}_{s_{\mathrm{eq}}}$, we split $\mathbf{Y}$ into $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$, where $\mathbf{Y}_{1}$ represents the results of tests corresponding to $X_{s_{\mathrm{eq}}}^{(i)}\neq 0$, and $\mathbf{Y}_{2}$ represents the remaining results corresponding to $X_{s_{\mathrm{eq}}}^{(i)}=0$. Due to the “OR” operation in the group testing model, $\mathbf{Y}_{1}$ is completely determined by the noise in its tests, which implies that $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ are conditionally independent given $\mathbf{X}_{s_{\mathrm{eq}}}$. Therefore, we have
$\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})$ $=\log\frac{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})}{\mathbb{P}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}})}$ (J.1)
$=\log\frac{\mathsf{P}(\mathbf{Y}_{1}|\mathbf{X}_{s_{\mathrm{dif}}})\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})}{\mathsf{P}(\mathbf{Y}_{1})\mathsf{P}(\mathbf{Y}_{2})}$ (J.2)
$=\log\frac{\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})}{\mathsf{P}(\mathbf{Y}_{2})},$ (J.3)
where (J.3) follows from $\mathsf{P}(\mathbf{Y}_{1}|\mathbf{X}_{s_{\mathrm{dif}}})=\mathsf{P}(\mathbf{Y}_{1})$ due to the fact that $\mathbf{Y}_{1}$ only depends on the noise. We let $n_{i}$ be the length of $\mathbf{Y}_{i}$ ($i=1,2$), which is deterministic given $\mathbf{X}_{s_{\mathrm{eq}}}$.
J.1 Proof of Lemma H.3 (Concentration Bound for $\frac{\ell}{k}\to\alpha\in[0,1]$)
@@
Concentration behavior of various quantities: We first consider the case $\alpha\in(0,1]$ and establish (H.10). Recall that we let $n_{i}$ be the length of $\mathbf{Y}_{i}$ for $i=1,2$. Let $M$ denote the number of tests corresponding to “$X_{s_{\mathrm{eq}}}^{(i)}=0,X_{s_{\mathrm{dif}}}^{(i)}\neq 0$”, [^11] Here, $M$ represents the random variable, while we will use lowercase letter $m$ for its specific value or actual realization; we will use a similar convention for other variables below, such as $V$ and $v$. meaning that $n_{2}-M$ is the number of tests corresponding to “$X_{s_{\mathrm{eq}}}^{(i)}=0,X_{s_{\mathrm{dif}}}^{(i)}=0$”, or equivalently $X^{(i)}_{s}=0$. This definition of $M$ coincides with $M_{s_{\mathrm{dif}}}$ in the notation introduced above Lemma B.1, but we omit the subscript to lighten notation, since we only consider one particular choice of $s_{\mathrm{dif}}$ here.
Note that $\frac{|s_{\mathrm{eq}}|}{k}=\frac{k-\ell}{k}=1-\frac{\ell}{k}$, and $\mathbf{Y}_{1}$ corresponds to the tests where “$X_{s_{\mathrm{eq}}}^{(i)}\neq 0$” so that we can write $n_{1}=W^{(s_{\mathrm{eq}})}$ in the notation of Lemma G.3. Thus, given $\delta>0$, (G.12) and (G.14) in Lemma G.3 give that with probability at least $1-2\exp\big{(}-2\nu^{-1}\delta^{2}n\big{)}$, we have
$n_{1}$ $=\Big{(}1-e^{-\nu(1-\frac{\ell}{k})(1+o(1))}+\delta^{\prime}\sqrt{1-\frac{\ell}{k}}\Big{)}n,\text{ for some }|\delta^{\prime}|<\delta,$ (J.4)
@@
$n_{2}=n-n_{1}=\Big{(}e^{-\nu(1-\frac{\ell}{k})(1+o(1))}-\delta^{\prime}\sqrt{1-\frac{\ell}{k}}\Big{)}n,\text{ for some }|\delta^{\prime}|<\delta.$ (J.5)
We note that (J.5) holds for all $\ell=1,\dotsc,k$. For the case of $\frac{\ell}{k}\to\alpha\in(0,1]$, we obtain from (J.5) that the event
$n_{2}=\big{(}e^{-\nu(1-\alpha)}+o(1)-\delta^{\prime}\big{)}n,\text{ for some }|\delta^{\prime}|<\delta.$ (J.6)
@@
Next, we seek to obtain a similar concentration result for $M$. Note that $n-(n_{2}-M)$ represents the number of tests where $X_{s}^{(i)}\neq 0$, so we can write $n-(n_{2}-M)=W^{(s)}$. Hence, (G.12) and (G.14) in Lemma G.3 yield that with probability at least $1-2\exp(-2\nu^{-1}\delta^{2}n)$, we have
$n-(n_{2}-M)=(1-e^{-\nu}+o(1)+\delta^{\prime\prime})n,\leavevmode\nobreak\ \text{for some }|\delta^{\prime\prime}|<\delta.$ (J.7)
Then, (J.6) and (J.7) imply
$M=(e^{-\nu(1-\alpha)}-e^{-\nu}+o(1)+\delta^{(3)})n,\leavevmode\nobreak\ \text{for some }|\delta^{(3)}|<2\delta.$ (J.8)
In the subsequent analysis, we will suppose that (J.6) and (J.8) hold, which is justified via a simple union bound for the multiple high-probability events.
@@
We take $t=n_{2}\delta$ to obtain that $|V-[M(1-\rho)+(n_{2}-M)\rho]|<n_{2}\delta$, or equivalently, $V=(1-2\rho)M+(\rho+\delta^{(4)})n_{2}$ for some $|\delta^{(4)}|<\delta$, with probability at least $1-2\exp\big{(}-2n_{2}\delta^{2}\big{)}$. Substituting the high-probability values of $M$ and $n_{2}$ in (J.6) and (J.8), when $\delta<\frac{1}{4}e^{-\nu(1-\alpha)}$, with probability at least $1-2\exp(-e^{-\nu(1-\alpha)}\delta^{2}n)$ we have
$V$ $=(1-2\rho)\big{(}e^{-\nu(1-\alpha)}-e^{-\nu}+o(1)+\delta^{(3)}\big{)}n+(\rho+\delta^{(4)})\big{(}e^{-\nu(1-\alpha)}+o(1)-\delta^{\prime}\big{)}n$ (J.10)
$=ne^{-\nu(1-\alpha)}\Big{(}(1-2\rho)(1-e^{-\nu\alpha})+\rho+o(1)+c_{1,\rho,\nu,\alpha}\delta\Big{)}$ (J.11)
$=ne^{-\nu(1-\alpha)}\big{(}\rho\star e^{-\nu\alpha}+o(1)+c_{1,\rho,\nu,\alpha}\delta\big{)},$ (J.12)
@@
where in (J.11) we introduce $c_{1,\rho,\nu,\alpha}$ that is bounded by an absolute constant depending on $(\rho,\nu,\alpha)$, [^12] Similarly, we will use $c_{2,\rho,\nu,\alpha}$, $c_{3,\rho,\nu,\alpha}$, etc. to denote constants depending on $(\rho,\nu,\alpha)$. in (J.12) we make the observation $(1-2\rho)(1-e^{-\nu\alpha})+\rho=\rho\star e^{-\nu\alpha}$ (recalling that $a\star b=ab+(1-a)(1-b)$).
@@
Recall that (J.3) gives $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})=\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})-\log\mathsf{P}(\mathbf{Y}_{2})$, and in both terms we implicitly condition on $\mathbf{X}_{s_{\mathrm{eq}}}$. To establish the concentration bound for $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})$, the main work is to bound these two terms separately.
@@
Bounding $\log\mathsf{P}(\mathbf{Y}_{2})$: To handle the term $\mathsf{P}(\mathbf{Y}_{2})$, consider a fixed realization $\mathbf{y}_{2}$ of $\mathbf{Y}_{2}$. This implies that the realization $v$ of $V$ is also fixed (since it is a function of $\mathbf{y}_{2}$), and we suppose that it satisfies the high-probability condition (J.12). By the symmetry of the test design with respect to re-ordering the test indices, all $\binom{n_{2}}{v}$ possibilities of $\mathbf{Y}_{2}$ having weight $v$ occur with equal probability, and hence, we have $\mathsf{P}(\mathbf{y}_{2})=\frac{\mathsf{P}(v)}{\binom{n_{2}}{v}}$. Then, applying [^13] This step may appear loose, but the idea is that if we were to consider all $v$ and take the dominant one, it would have a “large” value of $\mathbb{P}(v)$ anyway, i.e., not decaying fast enough to impact the final result. $\mathsf{P}(v)\leq 1$ and substituting (J.6) and (J.12), we obtain
$\mathsf{P}(\mathbf{y}_{2})\leq\frac{1}{\binom{ne^{-\nu(1-\alpha)}(1+o(1)+c_{2,\rho,\nu,\alpha}\delta)}{ne^{-\nu(1-\alpha)}(e^{-\alpha\nu}\star\rho+o(1)+c_{1,\rho,\nu,\alpha}\delta)}}.$ (J.13)
@@
From Lemma A.2, we have $\binom{N}{\beta N}\geq\frac{\exp(NH_{2}(\beta))}{\sqrt{2N}}=\exp(N[H_{2}(\beta)+o(1)])$ for constant $\beta\in(0,1)$, which implies that if $\delta$ is smaller than some constant depending on $(\alpha,\nu,\rho)$ such that $\frac{e^{-\alpha\nu}\star\rho+o(1)+c_{1,\rho,\nu,\alpha}\delta}{1+o(1)+c_{2,\rho,\nu,\alpha}\delta}$ is sufficiently close to $e^{-\alpha\nu}\star\rho$, we can upper bound the right-hand side of (J.13) by $\exp\big{(}-[1+o(1)+c_{3,\rho,\nu,\alpha}\delta]ne^{-\nu(1-\alpha)}H_{2}(e^{-\alpha\nu}\star\rho)\big{)}$ for some constant $c_{3,\rho,\nu,\alpha}$. Since this holds for any realization of $\mathbf{y}_{2}$ whose corresponding $v$ value satisfies (J.12), we conclude that
$\mathsf{P}(\mathbf{Y}_{2})\leq\exp\Big{(}-[1+o(1)+c_{3,\rho,\nu,\alpha}\delta]ne^{-\nu(1-\alpha)}H_{2}(e^{-\alpha\nu}\star\rho)\Big{)}$ (J.14)
@@
$\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})=\log\prod_{i=1}^{n_{2}}\mathsf{P}(Y_{2}^{(i)}|X_{s_{\mathrm{dif}}}^{(i)})=\sum_{i=1}^{n_{2}}\log\mathsf{P}(Y_{2}^{(i)}|X_{s_{\mathrm{dif}}}^{(i)}),$ (J.15)
where without loss of generality we index the tests in $\mathbf{Y}_{2}$ from $1$ to $n_{2}$. Note that (J.15) indicates that $\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})$ is a sum of $n_{2}$ independent random variables, which have a common expected value $-H_{2}(\rho)$ and are bounded by $|\log\rho|$ (since for any $\rho\in\big{(}0,\frac{1}{2}\big{)}$ we have $|\rho\log\rho+(1-\rho)\log(1-\rho)|\leq\rho|\log\rho|+(1-\rho)|\log\rho|=|\log\rho|$), so using Hoeffding’s inequality [12, Thm. 2.8] with respect to the noise randomness gives the following for any $t>0$:
$\mathsf{P}\Big{(}\big{|}\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})+n_{2}H_{2}(\rho)\big{|}\geq t\Big{)}\leq 2\exp\Big{(}-\frac{2t^{2}}{n_{2}\log^{2}\rho}\Big{)}.$ (J.16)
Moreover, we use (J.16) and let $t=\delta n_{2}H_{2}(\rho)$ to obtain that $|\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})+n_{2}H_{2}(\rho)|\leq\delta n_{2}H_{2}(\rho)$ with probability at least $1-2\exp\big{(}-2\delta^{2}(\log\rho)^{-2}(H_{2}(\rho))^{2}n_{2}\big{)}$. Further substituting (J.6), when $\delta<\frac{1}{4}e^{-\nu(1-\alpha)}$, we obtain that
$\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})=-ne^{-\nu(1-\alpha)}H_{2}(\rho)\big{(}1+c_{4,\rho,\nu,\alpha}\delta\big{)}$ (J.17)
@@
Putting the pieces together: We have shown that (J.14) and (J.17) hold with probability at least $1-\exp\big{(}-C_{\rho,\nu,\alpha}\delta^{2}n\big{)}$ for some $C_{\rho,\nu,\alpha}>0$ depending on $(\rho,\nu,\alpha)$ only (here we suppose that $\delta$ is given and $n$ is large enough, so that any leading constant before $\exp(-C_{\rho,\nu,\alpha}\delta^{2}n)$ can be absorbed into the exponent). Now we start with (J.3) and proceed as
$\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})=\log\mathsf{P}(\mathbf{Y}_{2}|\mathbf{X}_{s_{\mathrm{dif}}})-\log\mathsf{P}(\mathbf{Y}_{2})$ (J.18)
$\geq-ne^{-\nu(1-\alpha)}H_{2}(\rho)(1+c_{4,\rho,\nu,\alpha}\delta)+[1+o(1)+c_{3,\rho,\nu,\alpha}\delta]ne^{-\nu(1-\alpha)}H_{2}(e^{-\alpha\nu}\star\rho)$ (J.19)
$\geq(1-o(1))I_{\ell}^{n}+\big{(}c_{3,\rho,\nu,\alpha}H_{2}(e^{-\alpha\nu}\star\rho)-c_{4,\rho,\nu,\alpha}H_{2}(\rho)\big{)}\delta ne^{-\nu(1-\alpha)}$ (J.20)
$\geq(1-o(1))I_{\ell}^{n}-C_{1,\rho,\nu,\alpha}n\delta$ (J.21)
$=\Big{(}1-o(1)-\frac{C_{1,\rho,\nu,\alpha}n\delta}{I_{\ell}^{n}}\Big{)}I_{\ell}^{n},$ (J.22)
where in (J.19) we substitute (J.14) and (J.17), in (J.20) we recall $I_{\ell}^{n}$ given in (H.9) in Lemma H.2, and in (J.21) $C_{1,\rho,\nu,\alpha}$ is a constant depending on $(\rho,\nu,\alpha)$. Thus, given any $\delta_{1}>0$, we can set the above $\delta$ as $\delta=\frac{I_{\ell}^{n}\delta_{1}}{2nC_{1,\rho,\nu,\alpha}}$, and we obtain $\imath^{n}(\mathbf{Y}|\mathbf{X}_{s_{\mathrm{eq}}},\mathbf{X}_{s_{\mathrm{dif}}})\geq(1-o(1)-\frac{\delta_{1}}{2})I_{\ell}^{n}$ with probability at least $1-4\exp\big{(}-\frac{C_{\rho,\nu,\alpha}^{\prime}(I_{\ell}^{n})^{2}\delta_{1}^{2}}{n}\big{)}$ for some $C^{\prime}_{\rho,\nu,\alpha}>0$ only depending on $(\rho,\nu,\alpha)$, as desired.